A body of weight 50N is hauled along a tough horizontal plane by a pull of 18N acting at an angle of 14° with the horizontal. Find the coefficient of friction


1
Expert's answer
2021-06-17T15:08:10-0400

"F_1= 18N"

"\text{let }\vec F \text{ resultant force}"

"\vec F= m \vec a"

"\vec a = 0 \text{ since the body is hauled}"

"\vec F=0"

"\vec F = \vec F_1+\vec F_{fr}+\vec N"

"\text{for vertical axis Y:}"

"\vec N-mg+\vec F_1\sin14\degree= 0"

"\vec N =mg - \vec F_1\sin14\degree= 50-18*\sin14\degree=45.66"

"\text{for horizontal axis X:}"

"\vec F_1 \cos14\degree-F_{fr}=0"

"F_{fr}=\vec F_1 \cos14\degree=18*\cos14\degree=17.47"


"\mu= \frac{F_{fr}}{N}=\frac{17.47}{45.66}=0.383"

"\text{Answer: } \mu= 0.383"